A sample of 400 observations will be taken from a process (an infinite population). The population proportion equals 0.8. The probability that the sample proportion will be greater than 0.83 is _____.
Answer #1

Solution

Given that,

=  [p( 1 - p ) / n] = [(0.8 * 0.2) / 400 ] = 0.02

P( > 0.83) = 1 - P( < 0.83)

= 1 - P(( - ) / < (0.83 - 0.8) / 0.02)

= 1 - P(z < 1.5)

= 1 - 0.9332

= 0.0668

Probability = 0.0668